OpenStudy (anonymous):

Find the absolute maximums and minimums of f on the given closed interval and state where these values occur. f(x) = sinx-cosx

OpenStudy (anonymous):

The closed interval is 0,pi

OpenStudy (anonymous):

I can't do it manually but bu graphing it on a ti-84 between 0 and pi I get max y of 1.4142136 @ x of 2.3561932 and min y of -1 at x of 0.

OpenStudy (anonymous):

Did you use a graphing calculator?

OpenStudy (anonymous):

yes also, you may recognize the 1.414 ar sqrt2

OpenStudy (anonymous):

Alright well thanks for helping. You've got the best response so far :)

OpenStudy (anonymous):

is that a calc 1 question?

OpenStudy (anonymous):

AP Calc AB yeah

OpenStudy (anonymous):

write as a single function of sine, then it will be much easier

OpenStudy (anonymous):

sinx = cosx ? ?

OpenStudy (phi):

take the derivative, set = 0 ,

OpenStudy (anonymous):

well taking the derivative of the original equation, 0 = cosx + sinx

OpenStudy (anonymous):

\[a\sin(x)+b\cos(x)=\sqrt{a^2+b^2}\sin(x+\theta)\] where \[\cos(\theta)=\frac{a}{\sqrt{a^2+b^2}}, \sin(\theta)=\frac{b}{\sqrt{a^2+b^2}}\]

OpenStudy (phi):

divide by cos x

OpenStudy (anonymous):

in your case \(a=1,b=-1\) and so it is \[\sqrt{2}\sin(x-\frac{\pi}{4})\]

OpenStudy (anonymous):

0=sinx/cosx = tanx

OpenStudy (anonymous):

then no calc needed you see the answer from your eyeballs

OpenStudy (phi):

0 = cosx + sinx 0= 1 + tan x

OpenStudy (phi):

true sat, but he's taking an AP calc test

OpenStudy (anonymous):

the largest \[\sqrt2\sin(x-\frac{\pi}{4})\] can be is \(\sqrt{2}\) and the min is \(-\sqrt{2}\)

OpenStudy (anonymous):

true, but isn't is more obvious by symmetry that the largest \(\sin(x)-\cos(x)\) it can be is if \(\sin(x)=-\cos(x)\) ?

OpenStudy (anonymous):

sat, I am a little confused. What did you do after you found the derivative.

OpenStudy (anonymous):

i never found the derivative i used the identity \[\sin(x)-\cos(x)=\sqrt2\sin(x-\frac{\pi}{4})\]

OpenStudy (anonymous):

the amplitude is \(\sqrt{2}\) so that is the max

OpenStudy (anonymous):

Ok i see the problem, and maybe it is because my teacher does it slightly diffferent. How did you get sqrt2sin(x-pi/4)

OpenStudy (anonymous):

i gotta run, but i wrote all the steps for finding it above

OpenStudy (anonymous):

I'll look above. Thanks for the help

OpenStudy (anonymous):

yw

OpenStudy (phi):

just in case you want to use calc: d (sinx-cosx) = cos x + sin x = 0 1 + tan x = 0 tan x = -1 x = 135º and 315º (or 3pi/4 and -pi/4 ) in the interval 0 to pi, we only have 3pi/4 sin 3pi/4 = sqr(2)/2 cos 3pi/4 = - sqr(2)/2 sin - cos = sqr(2) to find the min, test the boundaries 0 and pi for the min value

OpenStudy (anonymous):

So that's where the sqrt2 came from. This looks more like what my teacher was showing me. I like this explanation better thank you.

OpenStudy (phi):

sat did not use calculus... more a trig approach. It is a technique people often use when solving differential equations.